Please help me solve. I'm stuck.
(y/y-1)^2=6(y/y-1)+7

Found 2 solutions by stanbon, xinxin:
Answer by stanbon(75887)

(y/y-1)^2=6(y/y-1)+7
------
Let y/(y-1) = w
------
Substitute to get:
w^2-6w-7 = 0
------
Factor:
(w-7)(w+1) = 0
x = 7 or w = -1
-------
Solve for "y":
y/(y-1) = 7 or y/(y-1) = -1
-----
y = 7y-7 or y = 1-y
-----
6y = 7 or 2y = 1
y = 7/6 or y = 1/2
======================
Cheers,
Stan H.

Answer by xinxin(76)

Let (y/y-1) = A, then:
A^2 = 6A +7
Solve for A, get A = 7 or A = -1.
When A = 7, y/y-1 = 7, y = 7/6; when A = -1, y/y-1 = -1, y = 1/2